100(1.11)^(x/6)=400
I can't figure out how to enter this equation into my scientific calculator to get an answer of 79.70 (answer rounded to the nearest hundredth)

Answer by RAY100(1637)

Remember loga^b = b*loga
.
100(1.11)^(x/6) = 400
.
1.11^(x/6) =4
.
log{1.11^(x/6) ) = log(4)
.
(x/6) * log(1.11) = log(4)
.
x/6 = log(4) /log(1.11) =13.2838
.
x=6*13.2838 = 79.7026
.
check
100(1.11)^(79.7026/6) = 400
400 = 400,,,ok